# Sparrow: Data-Efficient Video-LLM with Text-to-Image Augmentation

### v4

@@ 5.2 Ablation on Different Data Compositions
Table 2: Results of different data compositions on the Video-MME benchmark. Our proposed scheme achieves an overall performance superior to other data mixes of the same amount (30K) and even more data (200K). The Zero-shot and 200K full data fine-tuning settings are listed for comparison.
@@
[…]
@@
Figure 7: Scaling performance with different data volumes and types on general video benchmarks. We consider four settings, i.e. fine-tuning with video data only, our hybrid data (video data mixed with our synthetic samples with a ratio of 2:1:212:12 : 1), TOPA data, and T3 data. By default, video data (the caption dataset and the instruction dataset) are mixed with a 1:$1:11$ : 1 sampling ratio. Training samples are measured in K, where 0 sample indicates zero-shot inference. Purple-shaded area indicates gains of our methods over the video data baseline.
Can synthetic data help models understand longer videos? Interestingly, in the training stage, we only utilize synthetic samples of long multimodal context instead of authentic long video samples. However, on the long video benchmark set, our proposed method still achieves a score that is 1.2 points higher than the full data training (as shown in the 4th row compared to 200K full data in Tab. 2). This result suggests that fine-tuning with a long multimodal context can enhance the comprehension of longer videos. In the following section, we will present additional results and discussions to illustrate this point further.
@@ 5.3 Examination of Key Properties
In this section, we further examine key properties of our proposed method, including general effectiveness, scaling performance, and effectiveness when adopted in long video understanding scenarios.
@@ 5.3.1 General Effectiveness and Scaling Performance
We further verify the proposed scheme’s effectiveness by evaluating our methods on another image-LLM of larger parameter size, i.e. MiniCPM-8B, across different benchmarks. Specifically, through scaling up with different volumes and types of data, we compare our methods against a pure video data baseline, as well as other relevant methods, including TOPA and T3. Both TOPA and T3 first translate vision information into text, such as captions and relations between objects. Then, synthetic text QAs are constructed to simulate video reasoning samples, aiming to transfer temporal reasoning capabilities from text to video. Note that since the original data format of TOPA is largely different from the current paradigm, we design a template to adapt them to the instruction data format (More details are available in Appendix A.2). The results are summarized in Fig. 7.
@@
Scaling performance. A clear issue with other methods is that they are more prone to performance saturation when scaling up the data budget. For instance, the baseline method uses 60K samples to improve the Video-MME benchmark by 3.1 points, while using 100K samples only achieves another 0.9 points of absolute gains. In contrast, our proposed scheme shows more stable and consistent improvements when scaling up the data volumes compared with other methods. This suggests that ensuring a more diverse distribution of instruction in video training is very important, or otherwise, the learning efficiency may downgrade. To avoid such circumstances, the proposed data augmentation method can be an effective approach.
Discussion: Can we scale up only with synthetic textual samples? An intriguing and highly relevant question is whether we can scale up the synthetic samples without using any real video samples. Since the text is more compact and less redundant than a whole video, training in this way is more economical. Unfortunately, the empirical results show that this is probably unfeasible. As shown in Fig. 7, scaling with synthetic text data (TOPA and T3) shows undesirable characteristics, i.e. this approach can easily reach the saturation point or even slightly downgrade. Other critical issues include the modality gap and special processing of videos in various domains (such as egocentric videos and movies). Besides, since text suffers inevitably from information loss when translated from videos, text data might be better used as a supplement to videos, which helps inject temporal reasoning language prior into the LLM backbones (similar to TOPA or T3) or mixing with video data as a regularization method (as our method does).
Figure 8: Long video understanding performance evaluated with a different number of frames.Video context” denotes the max number of video frames used in training, i.e. 24 frames; “LLM context” denotes the context window of the LLM backbone. Purple-shaded area indicates performance gains of our proposed method over the video data baseline.
@@ 5.3.2 Long Video Understanding Performance
We adopt tailored benchmarks to evaluate long video understanding capabilities, including LongVideoBench [49], MLVU-M [50], and the long video set of Video-MME, and report the performance on evaluation sets for the former two benchmarks. Our study focuses on two aspects: (1) performance improvement in terms of long video understanding compared with the video fine-tuning baseline and (2) frame number (multimodal context) generalization ability in the inference stage.
Performance improvement w.r.t. training sample size. As shown in Tab. 3, we observe consistent improvements in long video understanding compared with the video fine-tuning baseline despite an absence of any long video training data. Notably, when using the same number of 100K training samples, our hybrid data scheme improves by 6.6 points over the baseline. We hypothesize this is because the reasoning patterns within the long text can also transfer to video understanding, which exempts the need to curate specific text samples translated from vision information [27, 28].
@@
Performance change w.r.t. frame number. We also investigate if our approach expands the context window. Larger context windows are usually beneficial since video-LLMs derived from large-context LLM backbones often accompany frame number generalization and benefit from inputting more video frames [49, 26]. However, we do not observe this trend. As shown in Fig. 8, the model fine-tuned with long synthetic samples still follows a similar pattern, that is, when performing inference beyond the video context of the training stage, the performance stays relatively stable and does not benefit from more frame input. And when the frame number exceeds the LLM context, the performance plunges to a low level. Thus, promising directions for further improvements may include continued pre-training to expand the LLM context window.
@@ 6 Conclusion
Appendix A More Implementation Details
A.1 Answer Judging
@@
To cope with these issues, we adopt a combination of exact matching and LLM matching for assessment. Specifically, we strip the prefixes such as “Answer:” from the prediction and try to use regular expression matching to find the option letter. When the exact matching scheme fails, we use an LLM (Llama-3.1-8B-Instruct [51]) to find an option closest to the model prediction. When the LLM matching fails, a placeholder outside of the available options (such as “Z”) is returned to denote a wrong answer. Our judging prompt for the LLM is modified from VLMEvalKit [52], as shown in Tab. 4.
A.2 Reproduction Details of Baseline Methods
@@
A real case of formatting the sample with the devised template is shown in Tab. 5.
@@
You are an AI assistant who will help me to match an answer with several options of a single-choice question.
@@
If the meaning of all options is significantly different from the answer, output Z. You should directly output a single uppercase character, such as A, B, C, D (if they are valid options) and Z, and nothing else. Here are two examples.

### v5

@@ 6.2 Ablation on Different Data Compositions
Table 2: Results of different data compositions on the Video-MME benchmark. Our proposed scheme achieves an overall performance superior to other data mixes of the same amount (30K) and even more data (200K).
@@
Examination of our design choices. As shown in Table 2, comparing the first three rows, we can find that when using the same amount of video samples, training only with ShareGemini is not as effective as using more diverse data compositions. Notably, under the same data budget, our proposed scheme (Row 4) achieves the best performance. Furthermore, compared to the full fine-tuning setting with 200K samples, our method achieves comparable performance using only 15% of the number of training samples. The training cost is also significantly reduced from 276.8 GPU hours to 33.6 GPU hours, yielding an 8.2$\times$ improvement in efficiency. These results underscore the importance of instruction diversity and validate the effectiveness of our proposed approach.
@@
Figure 7: Scaling performance with different data volumes and types on general video benchmarks. Specifically, we scale up the volume of training data until the overall performance gain saturates (empirically set as 1 point). Four settings are considered, i.e., fine-tuning with video data only, our hybrid data (video data mixed with our synthetic samples with a ratio of $2:1$), TOPA data, and T3 data. By default, video data (the caption dataset and the instruction dataset) are mixed with a $1:1$ sampling ratio. Training samples are measured in K, where 0 sample indicates zero-shot inference. Purple-shaded area indicates gains of our methods over the video data baseline.
Can synthetic data help models understand longer videos? Interestingly, in the training stage, we only utilize synthetic samples of long multimodal context instead of authentic long video samples. However, on the long video benchmark set, our proposed method still achieves a score that is 1.2 points higher than the full data training (as shown in the 4th row compared to 200K full data in Table 2). This result suggests that fine-tuning with a long multimodal context can enhance the comprehension of longer videos. In the following section, we will present additional results and discussions to illustrate this point further.
@@ 6.3 Examination of Key Properties
In this section, we further examine the key properties of our proposed method, including its general effectiveness, scaling performance, and effectiveness when applied to long video understanding scenarios.
@@ 6.3.1 General Effectiveness and Scaling Performance
We further verify the proposed scheme’s effectiveness by evaluating our methods on another image-LLM of larger parameter size, i.e., MiniCPM-8B, across different benchmarks. Through scaling up with different volumes and types of data, we compare our methods against a pure video data baseline, as well as other relevant methods, including TOPA and T3. Both TOPA and T3 first translate vision information into text, such as captions and relations between objects. Then, synthetic text QAs are constructed to simulate video reasoning samples, aiming to transfer temporal reasoning capabilities from text to video. Note that since the original data format of TOPA is largely different from the current paradigm, we design a template to adapt the samples to the instruction data format (More details are available in Appendix B). The results are summarized in Fig. 7.
@@
Scaling performance. A notable limitation of other methods is that they are more prone to performance saturation when scaling up the data budget. For instance, the baseline method uses 60K samples to improve the Video-MME benchmark by 3.1 points, while using 100K samples only achieves another 0.9 points of absolute gains. In contrast, our proposed scheme shows more stable and consistent improvements when scaling up the data volumes compared with other methods. This underscores the importance of maintaining instruction diversity in video-language training, as insufficient diversity can lead to reduced learning efficiency. In such cases, our data augmentation strategy facilitates a more diverse instructional distribution.
Discussion: Can we scale up only with synthetic textual samples? An intriguing and highly relevant question is whether we can scale up the synthetic samples without using any real video samples. Since the text is more compact and less redundant than a whole video, training in this way is more economical. Unfortunately, the empirical results show that this is probably unfeasible. As shown in Fig. 7, scaling with synthetic text data (TOPA and T3) shows undesirable characteristics, i.e., this approach can easily reach the saturation point or even slightly downgrade. Other critical issues include the modality gap and special processing of videos in various domains (such as egocentric videos and movies). Besides, since text suffers inevitably from information loss when translated from videos, text data might be better used as a supplement to videos, which helps inject temporal reasoning language prior into the LLM backbones (similar to TOPA or T3) or mixing with video data as a regularization method (as our method does).
Figure 8: Long video understanding performance evaluated with a different number of frames.Video context” denotes the max number of video frames used in training, i.e., 24 frames; “LLM context” denotes the context window of the LLM backbone. Purple-shaded area indicates performance gains of our proposed method over the video data baseline.
@@ 6.3.2 Long Video Understanding Performance
We adopt tailored benchmarks to evaluate long video understanding capabilities, including LongVideoBench [62], MLVU-M [63], and the long video set of Video-MME, and report the performance on evaluation sets for the former two benchmarks. Our study focuses on two aspects: (1) performance improvement in terms of long video understanding compared with the video fine-tuning baseline, and (2) frame number (multimodal context) generalization ability in the inference stage.
Performance change w.r.t. training configurations. As presented in Table 3, our method consistently outperforms the video-only fine-tuning baseline in long video understanding across varying sample sizes, despite the absence of any long video training data. Remarkably, with the same number of 100K training samples, our hybrid data strategy achieves a performance gain of 6.6 points over the baseline. We attribute this improvement to the transferable reasoning patterns embedded in long-form textual data, which may enhance the model’s temporal comprehension capabilities without requiring vision-derived textual supervision [41, 42].
We also examine whether adopting a denser sampling scheme can improve long video understanding performance, as this approach can also expand the context window. Specifically, we double the sampling frame limit (i.e., from 24 to 48) for authentic videos and find no improvements in long video understanding performance. This is because the training dataset predominantly contains short videos, and the dynamics within each video are relatively small. In this case, sampling more frames does not introduce more information but instead can bring redundancy, and thus, achieve no improvements in long video understanding.
@@
Performance change w.r.t. frame number. We also investigate whether our approach expands the context window. Larger context windows are usually beneficial since video-LLMs derived from large-context LLM backbones often accompany frame number generalization and benefit from inputting more video frames [62, 26]. However, we do not observe this trend. As shown in Fig. 8, the model fine-tuned with long synthetic samples still follows a similar pattern, that is, when performing inference beyond the video context of the training stage, the performance stays relatively stable and does not benefit from more frame input. And when the frame number exceeds the LLM context, the performance plunges to a low level. Thus, promising directions for further improvements may include continued pre-training to expand the LLM context window.
Figure 9: Qualitative results of OCR-related capabilities. The video features the process of dunking by a basketball player. Three settings are compared, i.e., zero-shot (direct inference with image-LLM), video data baseline (trained with 100K video samples), and our Sparrow scheme (trained with 66K video samples and 34K synthetic samples).
Figure 10: Model performance across various task categories on Video-MME. The Perception and Cognition categories are highlighted in the outer circle. Performance gains exceeding 5 points over the video data baseline are marked in green.
6.4 Fine-grained Performance Analysis
In this section, we present a more in-depth analysis of model performance. By examining qualitative results and task-wise performance breakdowns, we provide intuitive insights into how our method influences model performance.
6.4.1 Qualitative Results
We present qualitative results on Video-MME in Fig. 9, with additional examples provided in Appendix C. As shown in the figure, among the three models compared, only our method successfully identifies the number on the back of the basketball player, demonstrating superior OCR capabilities in video-based understanding. Further analyses on this aspect are discussed in the following section.
6.4.2 Performance Breakdown
To understand how our method influences model capabilities, we break down the performance across different tasks in Video-MME [38]. Three broad categories of specific capabilities involved in video understanding are considered:
Perception includes Counting Problems, OCR Problems, Object Recognition, Action Recognition, Attribute Perception, Spatial Perception, and Temporal Perception.
Cognition includes Temporal Reasoning, Spatial Reasoning, Action Reasoning, and Object Reasoning.
Information Synopsis. This category of questions emphasizes a model’s ability to comprehend the overall narrative or theme of a video rather than fine-grained details.
As shown in Fig. 10, beyond overall gains, the incorporation of synthetic data yields notable improvements in OCR-related tasks. In particular, the model trained with augmented synthetic samples outperforms the one trained on 100K real video samples by 10.4% in OCR problems. This notable gain aligns with expectations, as synthetic data enables the model to more effectively recognize and interpret textual elements embedded in text-intensive frames, thereby enhancing its ability to extract relevant information. Interestingly, a significant improvement is also observed in temporal reasoning, with a 7.9% performance gain. This observation suggests that the ability to comprehend textual context may partially transfer to the modeling of temporal relationships across multiple video frames.
@@ 7 Conclusion
Appendix A Answer Judging
@@
To cope with these issues, we adopt a combination of exact matching and LLM matching for assessment. Specifically, we strip the prefixes such as “Answer:” from the prediction and try to use regular expression matching to find the option letter. When the exact matching scheme fails, we use an LLM (Llama-3.1-8B-Instruct [64]) to find an option closest to the model prediction. When the LLM matching fails, a placeholder outside of the available options (such as “Z”) is returned to denote a wrong answer. Our judging prompt for the LLM is modified from VLMEvalKit [65], as shown in Table 4.
Appendix B Reproduction Details of Baseline Methods
@@
A real case of formatting the sample with the devised template is shown in Table 5.
@@
You are an AI assistant who will help me match an answer with several options of a single-choice question.
@@
If the meaning of all options is significantly different from the answer, output Z. You should directly output a single uppercase character, such as A, B, C, D (if they are valid options), and Z, and nothing else. Here are two examples.
@@ E. To fix the length of the wooden plank
Appendix C More Qualitative Results
Figure 11: A qualitative example from Video-MME. The video depicts humanoid robots performing a sequence of actions. Our model demonstrates a clear understanding of the video content by accurately detecting both motion patterns and the presence of two distinct robots.
Figure 12: A qualitative example from Video-MME. The video presents a tutorial on watercolor painting. In the zero-shot setting, the model exhibits an incorrect reasoning process, ultimately producing a conclusion that contradicts its own reasoning chain. In contrast, only our model demonstrates strong temporal reasoning abilities, successfully arriving at the correct answer.
In this section, we present more qualitative results, as illustrated in Fig. 11 and Fig. 12.